# On the Growth of Mistakes in Differentially Private Online Learning: A Lower Bound Perspective

### v1

@@ Abstract
In this paper, we provide lower bounds for Differentially Private (DP) Online Learning algorithms. Our result shows that, for a broad class of $\left({\varepsilon,\delta}\right)$-DP online algorithms, for $T$ such that $\log T\leqslant O\left(1/\delta\right)$, the expected number of mistakes incurred by the algorithm grows as $\Omega\left({\log\frac{T}{\delta}}\right)$. This matches the upper bound obtained by Golowich and Livni [19] and is in contrast to non-private online learning where the number of mistakes is independent of $T$. To the best of our knowledge, our work is the first result towards settling lower bounds for DP–Online learning and partially addresses the open question in Sanyal and Ramponi [27].
@@ 1 Introduction
There exists a hypothesis class $\mathcal{H}$ with $\mathrm{Ldim}\left({\mathcal{H}}\right)=1$ (see Definition 2), such that for any $\varepsilon,\delta>0,T\leqslant\exp(1/(16\delta))$, and any online learner $\mathcal{A}$ that is $(\varepsilon,\delta)$-DP and $\delta$-concentrated, there is an adversary, such that
$\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({\frac{\log T/\delta}{\varepsilon}}\right),$ (1.1)
where $\widetilde{\Omega}$ hides logarithmic factors in $\varepsilon$. For $T>\exp(1/(16\delta))$, $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({1/\delta}\right)$.
While the above result uses a hypothesis class of Littlestone dimension one, our main result in Section 4.1 also holds for any hypothesis class, even with Littlestone dimension greater than one. Utilising the $\mathrm{Point}_{N}$ hypothesis class (see Definition 4) in Section 1, we demonstrate that the minimum number of mistakes a DP online learner must make is bounded below by a term that increases logarithmically with the time horizon $T$. This holds if the learning algorithm is concentrated at least when $T$ is less than or equal to $\exp\left({1/(16\delta)}\right)$. This contrasts with non-private online learning, where the number of mistakes does not increase with $T$ in hypothesis classes with bounded Littlestone dimension, even if the learner is concentrated. [^1] All deterministic algorithms are $\operatorname{0-concentrated}$ by definition. Our result also shows that the analysis of the algorithm of Golowich and Livni [19], which shows an upper bound of $\Omega\left({\log\left({T/\delta}\right)}\right)$ for DP-SOA is tight as long as $T\leqslant\exp\left({1/(16\delta)}\right)$. However, as illustrated in Figure 1, this is not a limitation as for larger $T$, since a simple Name & Shame algorithm incurs lesser mistake than DP-SOA albeit at vacuous privacy levels (see discussion after Theorem 1).
@@ 2.1 Online Learning
Let $\mathcal{H}\subseteq\left\{0,1\right\}^{\mathcal{X}}$ be a hypothesis class of functions from $\mathcal{X}$ to $\left\{0,1\right\}$. The game between a learner $\mathcal{A}$ and adversary $\mathcal{B}$ is played as follows: adversary $\mathcal{B}$ picks $f^{\ast}\in\mathcal{H}$ and a sequence $x_{1},\ldots,x_{T}\in\mathcal{X}$ for $t=1,\ldots,T$: learner $\mathcal{A}$ outputs a current prediction $\hat{f}_{t}\in\left\{0,1\right\}^{\mathcal{X}}$ (see Definition 3) $\mathcal{A}$ receives $(x_{t},f^{\ast}(x_{t}))$ let $M_{\mathcal{A}}=\sum_{t=1}^{T}{\mathbb{I}}\left\{{\hat{f}_{t}(x_{t})\neq f^{\ast}(x_{t})}\right\}$
@@
[…]
@@ 2.2 Differential Privacy
Privacy of Prediction or Privacy of Predictor The first complexity arises from what the privacy adversary observes when altering an input, termed as its view. Since $\mathcal{A}$ provides an output hypothesis $\hat{f}_{t}\in\left\{{0,1}\right\}^{\mathcal{X}}$ at every time step $t\in[T]$ as shown in Definition 1, the adversary’s view could encompass the entire list of predictors. Our work, like Golowich and Livni [19], focuses on this scenario, where the output set is $S\subseteq\left\{0,1\right\}^{\mathcal{X}\times T}$. Nevertheless, certain studies restrict the adversary’s view to only the predictions, excluding the predictors themselves [13, 26]. In this setting, it is also important to assume that the adversary only observes the predictions on the inputs that it did not change [24, 22]; thus, they have $S\subseteq\left\{{0,1}\right\}^{\left({T-1}\right)}$.
@@ 3 Related work
Understanding which hypothesis classes can be privately learned is an area of vibrant research and was started in the context of Valiant’s PAC learning model [28]. A hypothesis class $\mathcal{H}$ is considered PAC-learnable if there exists an algorithm $\mathcal{A}$, which can utilize a polynomial-sized [^2] The term ’polynomial-sized’ refers to a sample size that is polynomial in the PAC parameters, including the error rate, confidence level, size of the hypothesis class, and the dimensionality of the input space., independent, and identically distributed (i.i.d.) sample $D$ from any data distribution to produce a hypothesis $h\in\mathcal{H}$ that achieves a low classification error with high probability on that distribution. In the context of DP-PAC learning, as defined by  Kasiviswanathan et al. [23], the learner $\mathcal{A}$ must also satisfy DP constraint with respect to the sample $D$. The overarching objective in this research domain is to find a clear criterion for the private learnability of hypothesis classes, analogous to the way learnability has been characterized in non-private settings—through the Vapnik-Chervonenkis (VC) dimension for offline learning [8] and the Littlestone dimension for online learning [25, 7].
@@ 4.1 Main Result
Let $\mathcal{H}$ be an arbitrary hypothesis class. Let $\varepsilon>0$, $\delta\leqslant\min(\varepsilon^{2},1/200)$ and $T\leqslant\exp(1/(16\delta))$. If $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and a $\delta$-concentrated online learner of $\mathcal{H}$, then there exists an adversary $\mathcal{B}$, such that
$\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({\frac{\log T/\delta}{\varepsilon}}\right),$ (4.2)
where $\widetilde{\Omega}$ contains logarithmic in $\varepsilon$ factors. For $T>\exp(1/(16\delta))$, $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({1/\delta}\right)$.
@@
Now, we compare result of Theorem 1 with known upper bounds in Figure 1. Recall that DP-SOA in Golowich and Livni [19] obtained an upper bound which increases logarithmically with the time horizon $T$Figure 1 shows that for $T\leqslant\exp(1/(16\delta))$, the dependency on $T$ is necessary, thereby showing the tightness of DP-SOA [19]. For larger $T$, the aforementioned Name & Shame actually outperforms DP-SOA, indicating that, for fixed $\varepsilon,\delta$, online algorithm always compromises privacy at very large $T$. However, even for smaller $T$, the plot shows that the number of mistake must grow with increasing $T$ until it matches the Name & Shame algorithm.
@@
Proof [of Theorem 1] Since $\mathcal{A}$ is $\operatorname{\delta-concentrated}$, there exists a distinguishing tuple $\left({f_{0},f_{1},x_{\mathrm{eq}},x_{\mathrm{dif}}}\right)$, such that
$\operatorname*{\mathbb{P}}\left[{\forall t\in[T],\ \mathcal{A}\left({\tau_{0}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=f_{0}(x_{\mathrm{dif}})}\right]\geqslant 1-\delta,$ (4.3)
where $\tau_{0}=\left({\left({x_{\mathrm{eq}},f_{0}(x_{\mathrm{eq}})}\right),\ldots,\left({x_{\mathrm{eq}},f_{0}(x_{\mathrm{eq}})}\right)}\right)$. WLOG assume that $f_{0}(x_{\mathrm{eq}})=f_{0}(x_{\mathrm{dif}})=0$. Note that by ignoring later rounds, we can assume $T\leqslant\exp(1/(16\delta))$. When $T>\exp(1/(16\delta))$, by simply only using the first $\exp(1/(16\delta))$ rounds, we obtain the required lower bound $\widetilde{\Omega}(1/\delta)$. Furthermore, let $k$ be the largest integer, such that $2^{k}-1\leqslant T$. Note that $\frac{1}{2}\log{T}\leqslant k\leqslant 2\log T\leqslant 1/(8\delta)$. For simplicity, we assume that $T=2^{k}-1$. We start with the case $\varepsilon=\varepsilon_{0}=\log\left({3/2}\right)$ and pick $f^{\ast}=f_{1}$ for the adversary.
Note that to show $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log T/\delta}\right)$, we can construct two adversaries, first achieving $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log 1/\delta}\right)$ (Case I) and the other $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log T}\right)$ (Case II). We now prove each of them.
Case I: We start with showing the adversary for the former bound, i.e., such that $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log 1/\delta}\right)$. The concentration assumption implies that for any $t\in[T]$, $\operatorname*{\mathbb{P}}\left({\mathcal{A}\left({\tau_{0}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)\leqslant\delta$. Therefore, if $\tau_{k}$ contains $k$ copies of the point $(x_{\mathrm{dif}},1)$, by applying DP property of $\mathcal{A}$ $k$ times, we obtain that for any $t\in[T]$, $\operatorname*{\mathbb{P}}\left({\mathcal{A}\left({\tau_{k}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)\leqslant\delta\left({\frac{\exp\left({\left({k+1}\right)\varepsilon_{0}}\right)-1}{\exp(\varepsilon_{0})-1}}\right)\leqslant k\delta\exp(k\varepsilon_{0})$. If $k=\frac{1}{4\varepsilon_{0}}\log(1/\delta)$, we derive that for any $t\in[T]$,
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{k})_{t}\left[{x_{\mathrm{dif}}}\right]=1)\leqslant k\delta\exp(k\varepsilon_{0})=\frac{\delta\log(1/\delta)}{4\delta^{1/4}\varepsilon_{0}}\leqslant\frac{1}{4}\delta^{1/4}\log(1/\delta)\leqslant\frac{1}{2},$ (4.4)
where the last inequality holds as long as $\delta<1/200$. This implies that for the sequence $\tau_{k}$, expected number of mistakes is $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]\geqslant\frac{1}{8\varepsilon_{0}}\log(1/\delta)$.
@@
    $p_{i}\coloneqq\operatorname*{\mathbb{P}}\left[{\forall t\in[1,l_{i}-1],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0}\right]\geqslant 1-4i\delta$,
@@
Assume that we obtain these sequences until $i=k\geqslant\frac{1}{2}\log T$. Then on the event $\left\{\forall t\in[1,l_{k}-1],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0\right\}$, $\mathcal{A}$ will make $k$ mistakes. Since $k\leqslant 1/(8\delta)$, from the second property above $p_{k}\geqslant 1/2$. This implies that for the sequence $\bm{c}^{(k)}$ we have $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]\geqslant p_{k}k\geqslant\frac{1}{4}\log T$.
We construct the sequences by induction, starting with $\bm{c}^{(0)}=\left({x_{\mathrm{eq}},\ldots,x_{\mathrm{eq}}}\right)$. For each $i$, $\bm{c}^{(i)}$ will differ from $\bm{c}^{(i+1)}$ at exactly one point. This allows us to use DP property of $\mathcal{A}$ in order to compare outputs on $\bm{c}^{(i)}$ and $\bm{c}^{(i+1)}$. From $\operatorname{\delta-concentrated}$ assumption, we can pick $l_{0}=1,r_{0}=T$ which gives $p_{0}=1$ and $q_{0}=\delta$ (we interpret $\operatorname*{\mathbb{P}}\left[{\forall t\in\emptyset\ldots}\right]=1$).
@@
which just account for the shift $l_{i}\to l_{i}+1$. Next, for $m_{i}=(l_{i}+r_{i})/2$ and for any $\bm{x}\in\left\{x_{\mathrm{eq}},x_{\mathrm{dif}}\right\}^{T}$, define
@@
If $\operatorname*{\mathbb{P}}\left({Q_{1}(\bm{c}^{(i+1)})}\right)\leqslant\operatorname*{\mathbb{P}}\left({Q_{2}(\bm{c}^{(i+1)})}\right)$ we set $l_{i+1}\coloneqq l_{i}+1,r_{i+1}\coloneqq m_{i}$, which gives $q_{i+1}=\operatorname*{\mathbb{P}}\left({Q_{1}(\bm{c}^{(i+1)})}\right)$ and $p_{i+1}=p^{\prime}_{i}\geqslant p_{i}-q_{i}$ Otherwise, we set $l_{i+1}=m_{i}+1,r_{i+1}=r_{i}$, with $q_{i+1}=\operatorname*{\mathbb{P}}\left({Q_{2}(\bm{c}^{(i+1)})}\right)$ and $p_{i+1}\geqslant p^{\prime}_{i}-q^{\prime}_{i}\geqslant p_{i}-2q_{i}$. Overall, with our choice of $l_{i+1},r_{i+1}$:
@@
By geometric sum properties, we have that $q_{i}\leqslant 2\delta$, and therefore, $p_{i}\geqslant 1-4i\delta$. Finally, note that $r_{i+1}-l_{i+1}=(r_{i}-l_{i})/2-1$. In the beginning we have $r_{0}-l_{0}=2^{k}-2$, therefore $r_{i}-l_{i}=2^{k-i}-2$. Thus, we can repeat this process $\frac{1}{2}\log T\leqslant k\leqslant 2\log T\leqslant 1/(8\delta)$ times. By construction, all inserted points for $\bm{c}^{(i)}$ $x_{\mathrm{dif}}$ lie on the prefix $[1,l_{i}-1]$.
In order to extend to other values of $\varepsilon$, we either insert more points per step (if $\varepsilon<\varepsilon_{0}$), or divide the segment into more parts (if $\varepsilon>\varepsilon_{0}$). We refer to Appendix D for the argument in these cases. This concludes the proof.

### v3

@@ Abstract
In this paper, we provide lower bounds for Differentially Private (DP) Online Learning algorithms. Our result shows that, for a broad class of $\left({\varepsilon,\delta}\right)$-DP online algorithms, for number of rounds $T$ such that $\log T\leqslant O\left(1/\delta\right)$, the expected number of mistakes incurred by the algorithm grows as $\Omega\left({\log T}\right)$. This matches the upper bound obtained by Golowich and Livni [21] and is in contrast to non-private online learning where the number of mistakes is independent of $T$. To the best of our knowledge, our work is the first result towards settling lower bounds for DP–Online learning and partially addresses the open question in Sanyal and Ramponi [29].
@@ 1 Introduction
There exists a hypothesis class $\mathcal{H}$ with $\mathrm{Ldim}\left({\mathcal{H}}\right)=1$ (see Definition 2), such that for any $\varepsilon,\delta>0,T\leqslant\exp(1/(32\delta))$, and any online learner $\mathcal{A}$ that is $(\varepsilon,\delta)$-DP and $0.1$-concentrated, there is an adversary, such that
$\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({\frac{\log T}{\varepsilon}}\right),$ (1.1)
where $\widetilde{\Omega}$ hides logarithmic factors in $\varepsilon$. For $T>\exp(1/(32\delta))$, $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({1/\delta}\right)$.
While the above result uses a hypothesis class of Littlestone dimension one, our main result in Section 4.1 also holds for any hypothesis class, even with Littlestone dimension greater than one. Utilising the $\mathrm{Point}_{N}$ hypothesis class (see Definition 4) in Section 1, we demonstrate that the minimum number of mistakes a DP online learner must make is bounded below by a term that increases logarithmically with the time horizon $T$. This holds if the learning algorithm is concentrated at least when $T$ is less than or equal to $\exp\left({1/(32\delta)}\right)$. This contrasts with non-private online learning, where the number of mistakes does not increase with $T$ in hypothesis classes with bounded Littlestone dimension, even if the learner is concentrated. [^1] All deterministic algorithms are $\operatorname{0-concentrated}$ by definition. Our result also shows that the analysis of the algorithm of Golowich and Livni [21], which shows an upper bound of $\Omega\left({\log\left({T/\delta}\right)}\right)$ for DP-SOA is tight as long as $T\leqslant\exp\left({1/(32\delta)}\right)$. However, as illustrated in Figure 1, this is not a limitation as for larger $T$, since a simple Name & Shame algorithm incurs lesser mistake than DP-SOA albeit at vacuous privacy levels (see discussion after Theorem 1).
@@ 2.1 Online Learning
Let $\mathcal{H}\subseteq\left\{0,1\right\}^{\mathcal{X}}$ be a hypothesis class of functions from $\mathcal{X}$ to $\left\{0,1\right\}$. The game between a learner $\mathcal{A}$ and adversary $\mathcal{B}$ is played as follows: {mdframed}[nobreak=true]
  • adversary $\mathcal{B}$ picks $f^{\ast}\in\mathcal{H}$ and a sequence $x_{1},\ldots,x_{T}\in\mathcal{X}$
  • for $t=1,\ldots,T$:
    • learner $\mathcal{A}$ outputs a current prediction $\hat{f}_{t}\in\left\{0,1\right\}^{\mathcal{X}}$ (see Definition 3)
    • $\mathcal{A}$ receives $(x_{t},f^{\ast}(x_{t}))$
  • let $M_{\mathcal{A}}=\sum_{t=1}^{T}{\mathbb{I}}\left\{{\hat{f}_{t}(x_{t})\neq f^{\ast}(x_{t})}\right\}$
@@
$\mathrm{Point}_{N}\coloneqq\left\{f^{(i)},i\in[N]\right\},\quad\text{where}\quad f^{(i)}(x)={\mathbb{I}}\left\{{i=x}\right\}.$ (2.2)
@@ 2.2 Differential Privacy
Privacy of Prediction or Privacy of Predictor The first complexity arises from what the privacy adversary observes when altering an input, termed as its view. Since $\mathcal{A}$ provides an output hypothesis $\hat{f}_{t}\in\left\{{0,1}\right\}^{\mathcal{X}}$ at every time step $t\in[T]$ as shown in Definition 1, the adversary’s view could encompass the entire list of predictors. Our work, like Golowich and Livni [21], focuses on this scenario, where the output set is $S\subseteq\left\{0,1\right\}^{\mathcal{X}\times T}$. Nevertheless, certain studies restrict the adversary’s view to only the predictions, excluding the predictors themselves [5, 14, 28]. In this setting, it is also important to assume that the adversary only observes the predictions on the inputs that it did not change [26, 24]; thus, they have $S\subseteq\left\{{0,1}\right\}^{\left({T-1}\right)}$.
@@ 3 Related work
Understanding which hypothesis classes can be privately learned is an area of vibrant research and was started in the context of Valiant’s PAC learning model [30]. A hypothesis class $\mathcal{H}$ is considered PAC-learnable if there exists an algorithm $\mathcal{A}$, which can utilize a polynomial-sized [^2] The term ‘polynomial-sized’ refers to a sample size that is polynomial in the PAC parameters, including the error rate, confidence level, size of the hypothesis class, and the dimensionality of the input space., independent, and identically distributed (i.i.d.) sample $D$ from any data distribution to produce a hypothesis $h\in\mathcal{H}$ that achieves a low classification error with high probability on that distribution. In the context of DP-PAC learning, as defined by  Kasiviswanathan et al. [25], the learner $\mathcal{A}$ must also satisfy DP constraint with respect to the sample $D$. The overarching objective in this research domain is to find a clear criterion for the private learnability of hypothesis classes, analogous to the way learnability has been characterized in non-private settings—through the Vapnik-Chervonenkis (VC) dimension for offline learning [8] and the Littlestone dimension for online learning [27, 7].
@@ 4.1 Main Result
Let $\mathcal{H}$ be an arbitrary hypothesis class. Let $\varepsilon>0$, $\delta\leqslant\varepsilon^{2}$ and $T\leqslant\exp(1/(32\delta))$. If, for some $\delta\leqslant\beta\leqslant 1/10$, $\mathcal{A}$ is a $\beta$-concentrated $(\varepsilon,\delta)$-DP online learner of $\mathcal{H}$, then there exists an adversary $\mathcal{B}$, such that
$\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({\frac{\log T/\beta}{\varepsilon}}\right),$ (4.2)
where $\widetilde{\Omega}$ contains logarithmic in $\varepsilon$ factors. For $T>\exp(1/(32\delta))$, $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\widetilde{\Omega}\left({1/\delta}\right)$.
@@
Now, we compare result of Theorem 1 with known upper bounds in Figure 1. Recall that DP-SOA in Golowich and Livni [21] obtained an upper bound which increases logarithmically with the time horizon $T$Figure 1 shows that for $T\leqslant\exp(1/(32\delta))$, the dependency on $T$ is necessary, thereby showing the tightness of DP-SOA [21]. For larger $T$, the aforementioned Name & Shame actually outperforms DP-SOA, indicating that, for fixed $\varepsilon,\delta$, online algorithm always compromises privacy at very large $T$. However, even for smaller $T$, the plot shows that the number of mistake must grow with increasing $T$ until it matches the Name & Shame algorithm.
@@
Proof [of Theorem 1] Since $\mathcal{A}$ is $0.1$-concentrated, there exists a distinguishing tuple $\left({f_{0},f_{1},x_{\mathrm{eq}},x_{\mathrm{dif}}}\right)$, such that
$\operatorname*{\mathbb{P}}\left[{\forall t\in[T],\ \mathcal{A}\left({\tau_{0}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=f_{0}(x_{\mathrm{dif}})}\right]\geqslant 0.9,$ (4.3)
where $\tau_{0}=\left({\left({x_{\mathrm{eq}},f_{0}(x_{\mathrm{eq}})}\right),\ldots,\left({x_{\mathrm{eq}},f_{0}(x_{\mathrm{eq}})}\right)}\right)$. WLOG assume that $f_{0}(x_{\mathrm{eq}})=f_{0}(x_{\mathrm{dif}})=0$. When $T>\exp(1/(32\delta))$, by simply only using the first $\exp(1/(32\delta))$ rounds, we obtain the required lower bound $\widetilde{\Omega}(1/\delta)$. In the remaining, we assume that $T\leqslant\exp(1/(32\delta))$. Furthermore, let $k$ be the largest integer, such that $2^{k}-1\leqslant T$. Note that $\frac{1}{2}\log{T}\leqslant k\leqslant 2\log T\leqslant 1/(16\delta)$. For simplicity, we also assume that $T=2^{k}-1$. We start with the case $\varepsilon=\varepsilon_{0}=\log\left({3/2}\right)$ and pick $f^{\ast}=f_{1}$ for the adversary.
Note that to show $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log T/\beta}\right)$, we can construct two adversaries, first achieving $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log 1/\beta}\right)$ (Case I) and the other $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log T}\right)$ (Case II). We now prove each of them.
Case I: We start with showing the adversary for the former bound, i.e., such that $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]=\Omega\left({\log 1/\beta}\right)$. The concentration assumption implies that for any $t\in[T]$, $\operatorname*{\mathbb{P}}\left({\mathcal{A}\left({\tau_{0}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)\leqslant\beta$. Therefore, if $\tau_{k}$ contains $k$ copies of the point $(x_{\mathrm{dif}},1)$, by applying DP property of $\mathcal{A}$ $k$ times, we obtain that for any $t\in[T]$, $\operatorname*{\mathbb{P}}\left({\mathcal{A}\left({\tau_{k}}\right)_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)\leqslant\beta\exp(k\varepsilon_{0})+\delta\left({\frac{\exp\left({k\varepsilon_{0}}\right)-1}{\exp(\varepsilon_{0})-1}}\right)\leqslant(k\delta+\beta)\exp(k\varepsilon_{0})$. If $k=\frac{1}{8\varepsilon_{0}}\log(1/\beta)$, we derive that for any $t\in[T]$,
$\operatorname*{\mathbb{P}}(\mathcal{A}(\tau_{k})_{t}\left[{x_{\mathrm{dif}}}\right]=1)\leqslant(k\delta+\beta)\exp(k\varepsilon_{0})\leqslant\frac{\delta\log(1/\delta)}{8\delta^{1/8}\varepsilon_{0}}\leqslant\frac{1}{3}\delta^{1/2}\log(1/\delta)\leqslant\frac{1}{2}.$ (4.4)
This implies that for the sequence $\tau_{k}$, expected number of mistakes is $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]\geqslant\frac{1}{16\varepsilon_{0}}\log(1/\beta)$.
@@
    $p_{i}\coloneqq\operatorname*{\mathbb{P}}\left[{\forall t\in[1,l_{i}-1],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0}\right]\geqslant 1/2-4i\delta$,
@@
Assume that we obtain these sequences until $i=k\geqslant\frac{1}{2}\log T$. Then on the event $\left\{\forall t\in[1,l_{k}-1],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0\right\}$, $\mathcal{A}$ will make $k$ mistakes. Since $k\leqslant 1/(16\delta)$, from the second property above $p_{k}\geqslant 1/4$. This implies that for the sequence $\bm{c}^{(k)}$ we have $\operatorname*{\mathbb{E}}\left[{M_{\mathcal{A}}}\right]\geqslant p_{k}k\geqslant\frac{1}{8}\log T$.
We construct the sequences by induction, starting with $\bm{c}^{(0)}=\left({x_{\mathrm{eq}},\ldots,x_{\mathrm{eq}}}\right)$. For each $i$, $\bm{c}^{(i)}$ will differ from $\bm{c}^{(i+1)}$ at exactly one point. This allows us to use DP property of $\mathcal{A}$ in order to compare outputs on $\bm{c}^{(i)}$ and $\bm{c}^{(i+1)}$. From $\operatorname{\delta-concentrated}$ assumption, we can pick $l_{0}=1,r_{0}=T$ which gives $p_{0}=1$ and $q_{0}=0.1$ (we interpret $\operatorname*{\mathbb{P}}\left[{\forall t\in\emptyset\ldots}\right]=1$).
@@
which just account for the shift $l_{i}\to l_{i}+1$. Note that, since the events are nested, we can compute $p^{\prime}_{i}-q^{\prime}_{i}=\operatorname*{\mathbb{P}}\left[{\forall t\in[1,r_{i}],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0}\right]=p_{i}-q_{i}$. Next, for $m_{i}=(l_{i}+r_{i})/2$ and for any $\bm{x}\in\left\{x_{\mathrm{eq}},x_{\mathrm{dif}}\right\}^{T}$, define
@@
If $\operatorname*{\mathbb{P}}\left({Q_{1}(\bm{c}^{(i+1)})}\right)\leqslant\operatorname*{\mathbb{P}}\left({Q_{2}(\bm{c}^{(i+1)})}\right)$ we set $l_{i+1}\coloneqq l_{i}+1,r_{i+1}\coloneqq m_{i}$, which gives $q_{i+1}=\operatorname*{\mathbb{P}}\left({Q_{1}(\bm{c}^{(i+1)})}\right)$ and $p_{i+1}=p^{\prime}_{i}\geqslant p_{i}-q_{i}$ When $\operatorname*{\mathbb{P}}\left({Q_{1}(\bm{c}^{(i+1)})}\right)>\operatorname*{\mathbb{P}}\left({Q_{2}(\bm{c}^{(i+1)})}\right)$, we set $l_{i+1}=m_{i}+1,r_{i+1}=r_{i}$, with $q_{i+1}=\operatorname*{\mathbb{P}}\left({Q_{2}(\bm{c}^{(i+1)})}\right)$. We can bound
$p_{i+1}$ $=\operatorname*{\mathbb{P}}\left({\forall t\in[1,m_{i}],\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=0}\right)$ (4.9)
$=\operatorname*{\mathbb{P}}\left({\forall t\in[1,l_{i}],\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=0}\right)$
$\quad-\operatorname*{\mathbb{P}}\left({\forall t\in[1,l_{i}],\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=0\textbf{ and }\exists t\in[l_{i}+1,m_{i}],\text{ s.t. }\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)$
$=p^{\prime}_{i}-\operatorname*{\mathbb{P}}\left({\forall t\in[1,l_{i}],\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=0\textbf{and }\exists t\in[l_{i}+1,m_{i}],\text{ s.t. }\mathcal{A}(\bm{c}^{(i+1)})_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)$
$\geqslant p^{\prime}_{i}-\frac{3}{2}\operatorname*{\mathbb{P}}\left({\forall t\in[1,l_{i}],\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=0\textbf{ and }\exists t\in[l_{i}+1,m_{i}],\text{ s.t. }\mathcal{A}(\bm{c}^{(i)})_{t}\left[{x_{\mathrm{dif}}}\right]=1}\right)$
$=p^{\prime}_{i}-\frac{3}{2}q^{\prime}_{i}=p_{i}-q_{i}-\frac{1}{2}q^{\prime}_{i}\geqslant p_{i}-\frac{3}{2}q_{i},$
where we used the DP property for the first inequality. Overall, we obtain with our choice of $l_{i+1},r_{i+1}$:
@@
By geometric sum properties, we have that $q_{i}\leqslant\frac{1}{10}(3/4)^{i}+2\delta$, and therefore, $p_{i}\geqslant 1/2-4i\delta$. Finally, note that $r_{i+1}-l_{i+1}=(r_{i}-l_{i})/2-1$. In the beginning we have $r_{0}-l_{0}=2^{k}-2$, therefore $r_{i}-l_{i}=2^{k-i}-2$. Thus, we can repeat this process $\frac{1}{2}\log T\leqslant k\leqslant 2\log T\leqslant 1/(16\delta)$ times. By construction, all inserted points $x_{\mathrm{dif}}$ for $\bm{c}^{(i)}$ lie on the prefix $[1,l_{i}-1]$. In order to extend to other values of $\varepsilon$, we either insert more points per step (if $\varepsilon<\varepsilon_{0}$), or divide the segment into more parts (if $\varepsilon>\varepsilon_{0}$). We refer to Appendix C for the argument in these cases. This concludes the proof.